About the GELLMU examples at this location

The examples here represent different ways of using generalized L^{A}TEX to generate SGML. In the case of the basic mode examples only the syntatic translator has been used. For each “regular GELLMU” example there is a full suite of the files generated by the didactic production system.

stem-name mode document type  command on chern syn. trans. function
ReadMe  regular article  gellmu  gellmu-trans
assgt2-custom regular article  gellmu  gellmu-trans
assgt2-example regular article  gellmu  gellmu-trans
gamma-iden  regular article  gellmu  gellmu-trans
ghack4  basic  hack  g2s(*)  gellmu-sgml
ghtml  basic  html  g2h  gellmu-html
wprod  basic  (x)html(+mathml) g2x  gellmu-xml

(*) By default the script g2s uses the syntactic translator function gellmu-trans, but for ghack4 the correct function is gellmu-sgml. Thus, for ghack4 the command line on chern should be

                                g2s ghack4 sgml